factor completely 2w^3x^4 -32w^3
Answer by user_dude2008(1862)

2w^3x^4 -32w^3 = 2w^3(x^4-16) = 2w^3(x^2-4)(x^2+4) = 2w^3(x+2)(x-2)(x^2+4)